OpenStudy (anonymous):

Describe the graphs of the functions f(x) = 2x - 1 and g(x) = -2x - 1. Compare and contrast the domain and range of f(x) and g(x).

OpenStudy (zzr0ck3r):

y=mx + b m = slope = rise over run of the line and b is where it crosses the y axis so if the slope is 2 = 2/1, and b = 3 then start at 3 on the y, then go up two right one, up two over one.....

OpenStudy (zzr0ck3r):

http://www.wolframalpha.com/input/?i=graph+y%3D2x%2B3+on+standard+axis look at the second graph to get an idea of what im talking about